# D-inverse constellations

### v1

@@ 5. Generalising constructions of Nambooripad and Lawson
There are likely to be many settings in which established results obtained using ordered groupoids may be more elegantly obtained using D-inverse constellations. It is therefore possible that inverse constellations will prove a useful tool in the study of both ordered groupoids and inverse semigroups. They also provide a perspective through which results for ordered groupoids may be able to be generalised, to constellations.
@@
The ESN Theorem was generalised significantly (although with category isomorphism replaced by category equivalence) when Nambooripad showed in [9] that the category of regular semigroups is equivalent to the category of ordered groupoids in which $D(\mathcal{Q})$ is a regular biordered set in a way compatible with its structure as an ordered groupoid. In light of Theorem 4.13, this fact could instead be stated in terms of D-inverse constellations rather than ordered groupoids.
@@
Suppose $S$ is a regular semigroup. As usual, let $V(a)\subseteq S$ denote the set of inverses of $a\in S$. As part of his “ESN theorem” for regular semigroups, Nambooripad showed that $P=\{(s,s^{\prime})\mid s\in S,s^{\prime}\in V(s)\}$ is an ordered groupoid, with $D((s,s^{\prime}))=(ss^{\prime},ss^{\prime})$, $R((s,s^{\prime}))=(s^{\prime}s,s^{\prime}s)$, and $(s,s^{\prime})\circ(t,t^{\prime})$ defined if and only if $s^{\prime}s=tt^{\prime}$ and equal to $(st,t^{\prime}s^{\prime})$, and $(s,s^{\prime})\leq(t,t^{\prime})$ means $s=ss^{\prime}t,s^{\prime}=t^{\prime}ss^{\prime}$ and $ss^{\prime}=ss^{\prime}tt^{\prime}=tt^{\prime}ss^{\prime}$.
@@
Now consider an involuted semigroup $S$, meaning a semigroup equipped with an involution satisfying $(st)^{*}=t^{*}s^{*}$ and $s^{**}=s$ for all $s,t\in S$. Let $E^{*}(S)=\{e\in S\mid e^{*}=e=e^{2}\}$, the set of projections in $S$. For all $s\in S$, Define $I^{*}(S)=\{s\in S\mid ss^{*}s=s\}$, the set of partial isometries of $S$. In Section $4.2$ of [8], Lawson shows that $I^{*}(S)$ forms an ordered groupoid in which the identities are the projections, $D(s)=ss^{*}$, $R(s)=s^{*}s$ and $s\circ t=st$ is defined if and only if $R(s)=D(t)$. Further, $s\leq t$ means $s=D(s)t$ with $D(s)\leq D(t)$ under the usual ordering of idempotents in $E^{*}(S)=\{ss^{*}\mid s\in I(S)\}$ given by $e\leq f$ whenever $e=ef(=fe)$. For $e\in E^{*}(S)$ and $s\in I^{*}(S)$ with $e\leq D(s)$ we define the restriction $e|s=es$, and dually for corestrictions.
@@
In what follows, we give a common generalisation of these facts, but using inverse constellations rather than ordered groupoids. The advantage of such an approach is that the order and restriction (and even range operation) need not be considered, only the constellation and domain (partial) operations.
@@
Let $S$ be a semigroup with $E$ a non-empty subset of $E(S)=\{e\in S\mid e^{2}=e\}$. In [2], an element $s\in S$ was said to be $E$-regular if there exists $t\in S$ for which $sts=s$ with $st,ts\in E$; in this case, there is even $u\in S$ such that $sus=s$ and $usu=u$, with $su,us\in E$ (simply let $u=tst$). Call such $u$ an $E$-inverse of $s$. (Semigroups in which every element has an $E$-inverse for some $E\subseteq E(S)$ were studied in [10].) In the case of an involuted semigroup $S$, if $s\in I^{*}(S)$, then $s^{*}$ is an $E^{*}(S)$-inverse of $s$ (since from $ss^{*}s=s$, it follows that $s^{*}ss^{*}=s^{*}$, and $ss^{*},s^{*}s\in E^{*}(S)$).
Let $R_{E}(S)$ be the set of $E$-regular elements of $S$, suppose $T\subseteq R_{E}(S)$ with $E\subseteq T$, and let
@@
For example, let $T=S$ be a semigroup in which $E(S)\neq\emptyset$, and then $E=E(S)$ is trivially $T$-normal since $s^{\prime}es\in E(S)$ if $s^{\prime}$ is an inverse of $s$ and $e$ is an idempotent for which $ess^{\prime}=e$. This recovers Nambooripad’s construction. Lawson’s example is the special case in which $S$ is an involuted semigroup, $E=E^{*}(S)$, the set of projections of $S$, and $T=I^{*}(S)$, the partial isometries of $S$: since $ss^{*}$ and $s^{*}s$ are projections, every partial isometry $s$ is $E^{*}(S)$-regular with inverse $s^{*}$. A set of idempotents $E\subseteq E(S)$ for a semigroup $S$ is reduced if for any $e,f\in E$ we have $ef=f$ if and only if $fe=f$. It is easy to see that if $E$ is reduced, then $E$-inverses, if they exist, are unique. In the case at hand, $E^{*}(S)$ is reduced. To see this, notice that for $e,f\in E^{*}(S)$, we have that $ef=f$ if and only if $fe=f^{*}e^{*}=(ef)^{*}=f^{*}=f$). Hence $s^{*}$ is the unique $E^{*}(S)$-inverse of $S$. The set $E^{*}(S)$ is also $T$-normal, since if $s\in T$ and $e\in E^{*}(S)$ with $e=es^{*}s$, then $s^{*}es\in E^{*}(S)$ as is easily seen.
@@
Suppose $S$ is a semigroup containing non-empty $E\subseteq E(S)$, with $E\subseteq T\subseteq R_{E}(S)$. Then $\mathcal{I}_{E}(T)=(I_{E}(T),\cdot,D)$ is an inverse constellation with $(s^{\prime},s)$ the D-inverse of $(s,s^{\prime})$ for all $(s,s^{\prime})\in I_{E}(S)$, if and only if $E$ is $T$-normal.
@@
Choose $(s,s^{\prime})\in I_{E}(T)$. Then $ss^{\prime}\in E\subseteq I_{E}(T)$, $(ss^{\prime},ss^{\prime})\in I_{E}(T)$, and $D((s,s^{\prime}))\cdot(s,s^{\prime})$ exists (since $(ss^{\prime})(ss^{\prime})=ss^{\prime}$ in $S$) and equals $(s,s^{\prime})$. If $(s,s^{\prime})=(e,e)\cdot(s,s^{\prime})=(es,s^{\prime}e)$ for some right identity $(e,e)$, then $ess^{\prime}=e$, $ss^{\prime}e=e$, $es=s$ and $s^{\prime}e=s^{\prime}$. Hence $ss^{\prime}=(es)(s^{\prime}e)=(ess^{\prime})e=ee=e$, showing uniqueness.
@@
Now if $(s,s^{\prime}),(t,t^{\prime}),(u,u^{\prime})\in I_{E}(T)$ with $(s,s^{\prime})\cdot(t,t^{\prime})=(st,t^{\prime}s^{\prime})$ and $(t,t^{\prime})\cdot(u,u^{\prime})=(tu,u^{\prime}t^{\prime})$ existing, then $s=stt^{\prime},tt^{\prime}s^{\prime}=s^{\prime},tuu^{\prime}=t$ and $uu^{\prime}t^{\prime}=t^{\prime}$, so $s(tuu^{\prime}t^{\prime})=stt^{\prime}=s$ and $(tuu^{\prime}t^{\prime})s^{\prime}=tt^{\prime}s^{\prime}=s^{\prime}$, so $(s,s^{\prime})\cdot((t,t^{\prime})\cdot(u,u^{\prime}))$ exists.
If $(s,s^{\prime})\cdot((t,t^{\prime})\cdot(u,u^{\prime}))$ exists then it equals $(s,s^{\prime})\cdot(tu,u^{\prime}t^{\prime})=(stu,u^{\prime}t^{\prime}s^{\prime})$ and we have $t=tuu^{\prime}$, $uu^{\prime}t^{\prime}=t^{\prime}$, $s=s(tu)(u^{\prime}t^{\prime})=stt^{\prime}$, and $s^{\prime}=(tu)(u^{\prime}t^{\prime})s^{\prime}=tt^{\prime}s^{\prime}$, so $(s,s^{\prime})\cdot(t,t^{\prime})=(st,t^{\prime}s)$ exists. Further, $(st)uu^{\prime}=st$ and $uu^{\prime}t^{\prime}s^{\prime}=t^{\prime}s^{\prime}$, and so $((s,s^{\prime})\cdot(t,t^{\prime}))\cdot(u,u^{\prime}))=(st,t^{\prime}s^{\prime})\cdot(u,u^{\prime})$ exists and equals $(stu,u^{\prime}t^{\prime}s^{\prime})=(s,s^{\prime})\cdot((t,t^{\prime})\cdot(u,u^{\prime}))$.
@@
It is easy to check normality. Finally, for all $s\in I_{E}(T)$, $(s,s^{\prime})\cdot(s^{\prime},s)$ exists (since $ss^{\prime}s=s$ and $s^{\prime}ss^{\prime}=s^{\prime}$ in $S$) and equals $(ss^{\prime},ss^{\prime})=D((s,s^{\prime}))$, and similarly $(s^{\prime},s)\cdot(s,s^{\prime})$ exists and equals $D((s^{\prime},s))$. We have completed the verification that $\mathcal{I}_{E}(T)$ is a D-inverse constellation.
Conversely, suppose that $\mathcal{I}_{E}(T)$ is a D-inverse constellation as in the statement of the theorem. Let $e\in E$ and suppose that $s^{\prime}$ is an $E$-inverse of $s$ for which $e=ess^{\prime}=ss^{\prime}e$. It follows that $(e,e)\cdot(s,s^{\prime})$ exists and equals $(es,s^{\prime}e)$, so $s^{\prime}e$ is an $E$-inverse of $es$. In particular then, $s^{\prime}es=s^{\prime}e(es)\in E$. Hence $E$ is $T$-normal. ∎
@@
As shown in [10], $E$-inverses in the semigroup $S$ are unique when they exist if and only if $E$ is pre-reduced, meaning that for all $e,f\in E$, $ef=f$ and $fe=e$ imply $e=f$, and $ef=e$ and $fe=f$ imply $e=f$.
@@
Suppose $S$ is a semigroup with $E\subseteq E(S)$, $T\subseteq R_{E}(S)$ with $E\subseteq T$. If $E$ is pre-reduced and $T$-normal, then $T$ is a $D$-inverse constellation, where we define $s\cdot t=st$ if and only if $stt^{\prime}=s$ and $tt^{\prime}s^{\prime}=s^{\prime}$, where $t^{\prime}$ is the unique $E$-inverse of $t$ and similarly for $s$, and the D-inverse of $s$ is its $E$-inverse $s^{\prime}$.
@@
The conditions on $S$ and $T$ ensure that any $s\in T$ has a unique $E$-inverse $s^{\prime}\in T$. Then, $(T,\cdot)\cong(I_{E}(T),\cdot)$ as partial algebras, via the isomorphism $s\leftrightarrow(s,s^{\prime})$, where $s^{\prime}$ is the unique $E$-inverse of $s\in S$, as is easily seen. That $T$ is D-inverse constellation now follows from Theorem 5.1. ∎
If $S$ is an involuted semigroup, recall that $I^{*}(S)$ is the set of partial isometries $T$ in $S$. As noted in the discussion prior to Theorem 5.1, $E^{*}(S)$ is reduced, hence pre-reduced, and is $I^{*}(S)$-normal. We can now apply Corollary 5.3 to recover Theorem 3 in Section $4.2$ in [8]: $I^{*}(S)$ is an inverse constellation, where we define $s\cdot t=st$ if and only if $stt^{*}=s$, hence it is an ordered groupoid in the way described in [8].
Note that in the situation of Corollary 5.3, if $s\cdot t$ exists for some $s,t\in T$, then $st\in T$ has $E$-inverse $t^{\prime}s^{\prime}\in T$; however, $t^{\prime}\cdot s^{\prime}$ may not be defined in the constellation $T$, so the law $(s\cdot t)^{\prime}=t^{\prime}\cdot s^{\prime}$ fails in general (since the left hand side may not exist).

### v2

@@ 5. Generalising constructions of Nambooripad and Lawson
There are likely to be many settings in which established results obtained using ordered groupoids may be more elegantly obtained using D-inverse constellations. It is therefore possible that D-inverse constellations will prove a useful tool in the study of both ordered groupoids and inverse semigroups. They also provide a perspective through which results for ordered groupoids may be able to be generalised to constellations.
@@
The ESN Theorem itself was generalised significantly (although with category isomorphism replaced by category equivalence) when Nambooripad showed in [10] that the category of regular semigroups is equivalent to the category of ordered groupoids in which $D(\mathcal{Q})$ is a regular biordered set in a way compatible with its structure as an ordered groupoid. In light of Theorem 4.14, this fact could instead be stated in terms of D-inverse constellations rather than ordered groupoids.
@@
Suppose that $S$ is a regular semigroup. As usual, let $V(a)\subseteq S$ denote the set of inverses of $a\in S$. As part of his “ESN theorem” for regular semigroups, Nambooripad showed that $P=\{(s,s^{\prime})\mid s\in S,s^{\prime}\in V(s)\}$ is an ordered groupoid, with $D((s,s^{\prime}))=(ss^{\prime},ss^{\prime})$, $R((s,s^{\prime}))=(s^{\prime}s,s^{\prime}s)$, and $(s,s^{\prime})\circ(t,t^{\prime})$ defined if and only if $s^{\prime}s=tt^{\prime}$ and equal to $(st,t^{\prime}s^{\prime})$, and $(s,s^{\prime})\leq(t,t^{\prime})$ means $s=ss^{\prime}t,s^{\prime}=t^{\prime}ss^{\prime}$ and $ss^{\prime}=ss^{\prime}tt^{\prime}=tt^{\prime}ss^{\prime}$.
@@
Now consider an involuted semigroup $S$, meaning a semigroup equipped with an involution, that is, a unary operation satisfying $(st)^{*}=t^{*}s^{*}$ and $s^{**}=s$ for all $s,t\in S$. Let $E^{*}(S)=\{e\in S\mid e^{*}=e=e^{2}\}$, the set of projections in $S$. Define $I^{*}(S)=\{s\in S\mid ss^{*}s=s\}$, the set of partial isometries of $S$. In Section $4.2$ of [9], Lawson shows that $I^{*}(S)$ forms an ordered groupoid in which the identities are the projections, $D(s)=ss^{*}$, $R(s)=s^{*}s$ and $s\circ t=st$ is defined if and only if $R(s)=D(t)$. Further, $s\leq t$ means $s=D(s)t$ with $D(s)\leq D(t)$ under the usual ordering of idempotents in $E^{*}(S)=\{ss^{*}\mid s\in I(S)\}$ given by $e\leq f$ whenever $e=ef(=fe)$. For $e\in E^{*}(S)$ and $s\in I^{*}(S)$ with $e\leq D(s)$ we define the restriction $e|s=es$, and dually for corestrictions.
@@
In what follows, we give a common generalisation of these facts, but using D-inverse constellations rather than ordered groupoids. The advantage of such an approach is that the order and restriction (and even range operation) need not be considered, only the constellation and domain (partial) operations.
@@
Let $S$ be a semigroup with $E$ a non-empty subset of $E(S)=\{e\in S\mid e^{2}=e\}$. In [2], an element $s\in S$ was said to be $E$-regular if there exists $t\in S$ for which $sts=s$ with $st,ts\in E$; if $s$ is $E$-regular, it follows that there is $u\in S$ such that $sus=s$ and $usu=u$, with $su,us\in E$ (simply let $u=tst$). Call such $u$ an $E$-inverse of $s$. (Semigroups in which every element has an $E$-inverse for some $E\subseteq E(S)$ were studied in [11].) In the case of an involuted semigroup $S$, if $s\in I^{*}(S)$, then $s^{*}$ is an $E^{*}(S)$-inverse of $s$ (since from $ss^{*}s=s$, it follows that $s^{*}ss^{*}=s^{*}$, and $ss^{*},s^{*}s\in E^{*}(S)$).
Let $R_{E}(S)$ be the set of $E$-regular elements of $S$, suppose that $T\subseteq R_{E}(S)$ with $E\subseteq T$, and let
@@
For example, let $T=S$ be a semigroup in which $E(S)\neq\emptyset$, and then $E=E(S)$ is trivially $T$-normal since $s^{\prime}es\in E(S)$ if $s^{\prime}$ is an inverse of $s$ and $e$ is an idempotent for which $ess^{\prime}=e$. This recovers Nambooripad’s construction. Lawson’s example is the special case in which $S$ is an involuted semigroup, $E=E^{*}(S)$, and $T=I^{*}(S)$, the partial isometries of $S$: since $ss^{*}$ and $s^{*}s$ are projections, every partial isometry $s$ is $E^{*}(S)$-regular with inverse $s^{*}$. A set of idempotents $E\subseteq E(S)$ for a semigroup $S$ is reduced if for any $e,f\in E$ we have $ef=f$ if and only if $fe=f$. It is easy to see that if $E$ is reduced, then $E$-inverses, if they exist, are unique. In the case at hand, $E^{*}(S)$ is reduced. To see this, notice that for $e,f\in E^{*}(S)$, we have that $ef=f$ if and only if $fe=f^{*}e^{*}=(ef)^{*}=f^{*}=f$. Hence $s^{*}$ is the unique $E^{*}(S)$-inverse of $S$. The set $E^{*}(S)$ is also $T$-normal, since if $s\in T$ and $e\in E^{*}(S)$ with $e=es^{*}s$, then $s^{*}es\in E^{*}(S)$, as is easily seen.
@@
Suppose that $S$ is a semigroup containing non-empty $E\subseteq E(S)$, with $E\subseteq T\subseteq R_{E}(S)$. Then $\mathcal{I}_{E}(T)=(I_{E}(T),\cdot,D)$ is a D-inverse constellation with $(s^{\prime},s)$ the D-inverse of $(s,s^{\prime})$ for all $(s,s^{\prime})\in I_{E}(S)$, if and only if $E$ is $T$-normal.
@@
Choose $(s,s^{\prime})\in I_{E}(T)$. Then $ss^{\prime}\in E\subseteq I_{E}(T)$, $(ss^{\prime},ss^{\prime})\in I_{E}(T)$, and $D((s,s^{\prime}))\cdot(s,s^{\prime})$ exists (since $(ss^{\prime})(ss^{\prime})=ss^{\prime}$ in $S$) and equals $(s,s^{\prime})$. If $(s,s^{\prime})=(e,e)\cdot(s,s^{\prime})=(es,s^{\prime}e)$ for some right identity $(e,e)$, then $^{\prime}=e, sse=ess^{\prime}$e=e, $es=s$ and $s^{\prime}e=s^{\prime}$. Hence $ss^{\prime}=(es)(s^{\prime}e)=(ess^{\prime})e=ee=e$, showing uniqueness.
@@
If $(s,s^{\prime}),(t,t^{\prime}),(u,u^{\prime})\in I_{E}(T)$ with $(s,s^{\prime})\cdot(t,t^{\prime})=(st,t^{\prime}s^{\prime})$ and $(t,t^{\prime})\cdot(u,u^{\prime})=(tu,u^{\prime}t^{\prime})$ both existing, then $s=stt^{\prime},tt^{\prime}s^{\prime}=s^{\prime},tuu^{\prime}=t$ and $uu^{\prime}t^{\prime}=t^{\prime}$, so $s(tuu^{\prime}t^{\prime})=stt^{\prime}=s$ and $(tuu^{\prime}t^{\prime})s^{\prime}=tt^{\prime}s^{\prime}=s^{\prime}$, giving that $(s,s^{\prime})\cdot((t,t^{\prime})\cdot(u,u^{\prime}))$ exists.
If $(s,s^{\prime})\cdot((t,t^{\prime})\cdot(u,u^{\prime}))$ exists then it equals $(s,s^{\prime})\cdot(tu,u^{\prime}t^{\prime})=(stu,u^{\prime}t^{\prime}s^{\prime})$ and we have $^{\prime}, uut=tuu^{\prime}t^{\prime}=t^{\prime}$, $s=s(tu)(u^{\prime}t^{\prime})=stt^{\prime}$, and $s^{\prime}=(tu)(u^{\prime}t^{\prime})s^{\prime}=tt^{\prime}s^{\prime}$, so $(s,s^{\prime})\cdot(t,t^{\prime})=(st,t^{\prime}s)$ exists. Further, $(st)uu^{\prime}=st$ and $uu^{\prime}t^{\prime}s^{\prime}=t^{\prime}s^{\prime}$, and so $((s,s^{\prime})\cdot(t,t^{\prime}))\cdot(u,u^{\prime}))=(st,t^{\prime}s^{\prime})\cdot(u,u^{\prime})$ exists and is equal to $(stu,u^{\prime}t^{\prime}s^{\prime})=(s,s^{\prime})\cdot((t,t^{\prime})\cdot(u,u^{\prime}))$.
@@
It is easy to check normality. Finally, for all $s\in I_{E}(T)$, we have $(s,s^{\prime})\cdot(s^{\prime},s)$ exists (since $ss^{\prime}s=s$ and $s^{\prime}ss^{\prime}=s^{\prime}$ in $S$) and equals $(ss^{\prime},ss^{\prime})=D((s,s^{\prime}))$, and similarly $(s^{\prime},s)\cdot(s,s^{\prime})$ exists and equals $D((s^{\prime},s))$. We have completed the verification that $\mathcal{I}_{E}(T)$ is a D-inverse constellation.
Conversely, suppose that $\mathcal{I}_{E}(T)$ is a D-inverse constellation as in the statement of the theorem. Let $e\in E$ and suppose that $s^{\prime}$ is an $E$-inverse of $s$ for which $e=ess^{\prime}=ss^{\prime}e$. It follows that $(e,e)\cdot(s,s^{\prime})$ exists and equals $(es,s^{\prime}e)$, so $s^{\prime}e$ is an $E$-inverse of $es$. In particular, $s^{\prime}es=s^{\prime}e(es)\in E$. Hence $E$ is $T$-normal. ∎
@@
As shown in [11], $E$-inverses in the semigroup $S$ are unique when they exist if and only if $E$ is pre-reduced, meaning that for all $\in E, ef=fe$f=f and $fe=e$ imply $e=f$, and $ef=e$ and $fe=f$ imply $e=f$.
@@
Suppose that $S$ is a semigroup with $E\subseteq E(S)$, $T\subseteq R_{E}(S)$ with $E\subseteq T$. If $E$ is pre-reduced and $T$-normal, then $T$ is a $D$-inverse constellation, where we define $s\cdot t=st$ if and only if $stt^{\prime}=s$ and $tt^{\prime}s^{\prime}=s^{\prime}$, where $t^{\prime}$ is the unique $E$-inverse of $t$ and similarly for $s$, and the D-inverse of $s$ is its $E$-inverse $s^{\prime}$.
@@
The conditions on $S$ and $T$ ensure that any $s\in T$ has a unique $E$-inverse $s^{\prime}\in T$. Then, $(T,\cdot)\cong(I_{E}(T),\cdot)$ as partial algebras, via the isomorphism $s\leftrightarrow(s,s^{\prime})$, where $s^{\prime}$ is the unique $E$-inverse of $s\in S$, as is easily seen. That $T$ is a D-inverse constellation now follows from Theorem 5.1. ∎
If $S$ is an involuted semigroup, recall that $I^{*}(S)$ is the set of partial isometries $T$ in $S$. As noted in the discussion prior to Theorem 5.1, $E^{*}(S)$ is reduced, hence pre-reduced, and is $I^{*}(S)$-normal. We can now apply Corollary 5.3 to recover Theorem 3 in Section $4.2$ in [9]: $I^{*}(S)$ is a D-inverse constellation, where we define $s\cdot t=st$ if and only if $stt^{*}=s$, hence it is an ordered groupoid in the way described in [9].
Note that in the situation of Corollary 5.3, if $s\cdot t$ exists for some $s,t\in T$, then $st\in T$ has $E$-inverse $t^{\prime}s^{\prime}\in T$; however, $t^{\prime}\cdot s^{\prime}$ may not be defined in the constellation $T$, so the law $(s\cdot t)^{\prime}=t^{\prime}\cdot s^{\prime}$ fails in general (since the right hand side may not exist).